# A note on real algebraic maps which are topologically special generic maps

### 21 Jan 2024 v3

@@ Abstract.
We investigate the bundle in Main Theorem 3 (1) and (2). Let the bundle denoted by $M_{{\rm Boundary},0}:=M_{{\rm B},0}$. First the bundle $M_{{\rm B},0}$ is a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this). We apply an argument used to show Proposition 1 in a revised style. More precisely, Proposition 2.1 of [39] is a corresponding result. It also refers to [32] there. We consider the restriction of the bundle $W_{0}-S(f_{0})$ over ${\rm Int}\ {\bar{N}}_{0}$ to ${\bar{N}}_{0}-N(\partial{\bar{N}}_{0})\subset{\rm Int}\ {\bar{N}}_{0}$. We remove the total space of this bundle obtained by the restriction from $W_{0}$. We have a smooth bundle $W_{1}$ over $\partial W_{f_{0}}$ whose fiber is homeomorphic to $D^{m-n+2}$. This is diffeomorphic to $D^{m-n+2}$ in the case $m-n+2\neq 4,5$. This is also thanks to Ehresmann’s fibration theorem (or a relative version of this). Hereafter we apply this implicitly in several scenes. The bundle $M_{{\rm Boundary},0}=M_{{\rm B},0}$ is regarded as a subbundle of the bundle $W_{1}$ over $\partial W_{f_{0}}$. The structure group of the bundle $M_{{\rm B},0}$ is restricted to the structure group of the (bundle canonically regarded as a) subbundle defined on the preimage of $\partial{\bar{N}}_{0}\times\{t\}$ for each $0<t\leq 1$. For each $0<t\leq 1$, it is a subbundle of the bundle obtained naturally as the subbundle of $W_{1}$ by restricting the total space to the preimage of $\partial{\bar{N}}_{0}\times\{t\}$ similarly. For each two obtained smooth bundles, fibers are regarded as $\partial D^{m-n+1}$ and $D^{m-n+1}$, respectively. The former bundle is also obtained by restricting the fiber to the boundary $\partial D^{m-n+1}$ of course. We consider an element ${\psi}_{t}$ of the structure group of the subbundle $M_{{\rm B},0,t}$ of $M_{{\rm B},0}$ defined for each $0<t\leq 1$ before. This element can be regarded as an element extended to an element preserving the distance between each point and the origin $0\in D^{m-n+1}$. This is also regarded as an element of the structure group of the subbundle $W_{1,t}$ of $W_{1}$ defined for each $0<t\leq 1$ before. For such an element, we can consider $\frac{1}{s}{\psi}_{t}(sv)$ for $v\in D^{m-n+1}$ which is not the origin and $0<s\leq 1$. Of course $v$, $sv$ and $\frac{1}{s}{\psi}_{t}(sv)$ are canonically regarded as vectors and hereafter we do not explain about this rule explicitly. Here $||\frac{1}{s}{\psi}_{t}(sv)||=||v||$. We consider the limit at $s=0$. For a fixed $v$, we have the value of the differential of ${\psi}_{t}$ at $v$, seen naturally as a tangent vector at the origin of ${D}^{m-n+1}$. Thus we naturally have an element regarded as a linear transformation of ”the unit disk $D^{m-n+1}$”. This means that the structure group reduces to the group isomorphic to the structure group of a linear bundle whose fiber is $D^{m-n+1}$. These arguments and the structure of the map around $S(f_{0})$ and $\partial{\bar{N}}_{0}$ complete the proof of Main Theorem 3 (1). We can also see that the bundle $M_{{\rm B},0}$ is trivial for the structure group. This completes the proof of Main Theorem 3 (2).
@@
By our arguments, we can also check Main Theorem 3 (3) easily. A result close to this is also presented in Theorem 1 (3) of [15]. ”$\overline{D^{\prime}}$” is seen as the ($m+1$)-dimensional bounded and compact manifold $W_{0}$ bounded by $M_{0}$ there. The topology of the ($m+1$)-dimensional manifold is not so precisely discussed there: only exposition on ”homotopy” is presented there.
@@
The next remark is related to structures of our new real algebraic maps.
Remark 2.
It seems that our new real algebraic maps are not special generic in general in the smooth category. It is important to know whether our maps are special generic or not. It is also difficult.
As another problem which is in some sense similar to this problem, it is also very difficult to know whether two homeomorphisms $\Phi$ and $\phi$ in Main Theorem 1 can be chosen as diffeomorphisms in general.
@@
[15] has no essential errors and we have also found several more rigorous arguments. These arguments are discussed in the present paper. We will revise the version and submit the refereed journal based on comments and our arguments and idea here. Note also that the content of [15] is independent of the content of the present paper essentially.
@@
This new study is based on studies on special generic maps and explicit construction of real algebraic functions and maps of the author. One of related events motivating the author to present this study is the conference ”Extension of the singularity theory” (https://www2.akita-nct.ac.jp/kasedou/workshop/rims2023/index.html)” and related discussions on singularity theory, differential geometry, and real algebraic geometry.
The author would like to thank Takahiro Yamamoto for asking some questions on the talk by the author, given on 2023/11/27, and related informal discussions at Kyushu University held on 2023/12/11, after the conference. He knows singularity theory and applications to geometry widely. Based on his knowledge and idea, he has first commented that projective spaces admit no special generic maps in considerable cases and that it is interesting: this is on the study [13] of the author. He also says that it is also important to consider so-called general fold maps and explicit classes for understanding projective spaces in geometric ways: fold maps are defined naturally generalizing the form of the last ($n$-th) component in the definition of a special generic map (see related pioneering studies [45, 47], studies [5, 6] on existence, [38], a pioneering studies on algebraic topology and differential topology of manifolds admitting fold maps, and [9] for related fundamental and advanced theory from singularity theory for example).
The author would like to thank Atsufumi Honda for asking questions and related informal discussions at Kyushu University after the conference on this theorem, which was an event independent of the previous discussion and held on 2023/12/8. It is also on non-existence of special generic maps on projective spaces and interesting problems from the viewpoint of differential geometry. As a researcher on singularity theory and applications to differential geometry, he says that Reeb’s theorem on Morse functions and special generic maps seem to be closely related to minimizing the total curvatures of curves, surfaces and Riemannian manifolds.
Last, in the informal discussions with Takahiro Yamamoto, the author also says that he is now studying differential geometry and real algebraic geometry further to understand manifolds in related categories, which are finer than the smooth category. We have also discussed the discussions with Atsufumi Honda. He has also commented that special generic maps obtain some nice global structures of manifolds.
In conclusion, we have enjoyed the discussions to present this study.

### 15 Feb 2024 v4

@@ Abstract.
We can present a simple and natural proof. However, we present a complicated proof to respect general cases. Let the bundle in Main Theorem 3 (1) and (2) denoted by $M_{{\rm Boundary},0}:=M_{{\rm B},0}$. First the bundle $M_{{\rm B},0}$ is a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this). We apply an argument used to show Proposition 1 in a revised style. More precisely, Proposition 2.1 of [43] is a corresponding result. It also refers to [36] there. We consider the restriction of the bundle $W_{0}-S(f_{0})$ over ${\rm Int}\ {\bar{N}}_{0}$ to ${\bar{N}}_{0}-N(\partial{\bar{N}}_{0})\subset{\rm Int}\ {\bar{N}}_{0}$. From $W_{0}$, we remove the total space of this bundle over ${\bar{N}}_{0}-N(\partial{\bar{N}}_{0})$. We have a smooth bundle $W_{1}$ over $\partial W_{f_{0}}$ whose fiber is homeomorphic to $D^{m-n+2}$. This fiber is diffeomorphic to $D^{m-n+2}$ in the case $m-n+2\neq 4,5$. This is also thanks to Ehresmann’s fibration theorem (or a relative version of this). The bundle $M_{{\rm Boundary},0}=M_{{\rm B},0}$ is regarded as a subbundle of the bundle $W_{1}$ over $\partial W_{f_{0}}$. The structure group of the bundle $M_{{\rm B},0}$ is naturally regarded as the structure group of the (bundle canonically regarded as a) subbundle defined on the preimage of $\partial{\bar{N}}_{0}\times\{t\}$ for each $0<t\leq 1$. For each $0<t\leq 1$, it is a subbundle of the following bundle: the bundle obtained naturally as the subbundle of $W_{1}$ by restricting the total space to the preimage of $\partial{\bar{N}}_{0}\times\{t\}$ similarly. For each two obtained smooth bundles for $0<t\leq 1$, fibers are regarded as a manifold homeomorphic to $\partial D^{m-n+1}$ and a manifold homeomorphic to $D^{m-n+1}$, respectively. The former bundle is also obtained by restricting the fiber to the boundary homeomorphic $\partial D^{m-n+1}$ of course. The manifold homeomorphic to $D^{m-n+1}$ is diffeomorphic to $D^{m-n+1}$ if $m-n+1\neq 4,5$ or the boundary is diffeomorphic to $\partial D^{m-n+1}=S^{m-n}$ with the constraint $m-n+1\neq 4$ of course. We consider an element ${\psi}_{t}$ of the structure group of the subbundle $M_{{\rm B},0,t}$ of $M_{{\rm B},0}$ defined for each $0<t\leq 1$ before. This element can be regarded as an element canonically extended to an element of the structure group of the subbundle $W_{1,t}$ of $W_{1}$ defined for each $0<t\leq 1$ before.
By these arguments, we may discuss the bundles and the structure groups in the following simple situation essentially. We can regard the structure group of the bundle $W_{1,t}$ as a group acting on the unit disk $D^{m-n+1}$ preserving the distance of each point and the origin $0\in D^{m-n+1}$. We explain about an element ${\psi}_{t}$ of the group. We can consider $\frac{1}{s}{\psi}_{t}(sv)$ for $v\in D^{m-n+1}-\{0\}$ and $0<s\leq 1$. Of course 00 here is the origin and $v$, $sv$ and $\frac{1}{s}{\psi}_{t}(sv)$ are canonically regarded as vectors. Here we may regard the relation $||\frac{1}{s}{\psi}_{t}(sv)||=||v||$ by considering the distance $||x||$ for $x\in D^{m-n+1}-\{0\}\subset{\mathbb{R}}^{m-n+1}$.
We consider the limit at $s=0$. For a fixed $v$, we have the value of the differential of ${\psi}_{t}$ at $v$, seen naturally as a tangent vector at the origin of ${D}^{m-n+1}$. Thus we naturally have an element regarded as a linear transformation of ”the unit disk $D^{m-n+1}$”. This means that the structure group reduces to the group isomorphic to the structure group of a linear bundle whose fiber is $D^{m-n+1}$.
These arguments and the structure of the map around $S(f_{0})$ and $\partial{\bar{N}}_{0}$ complete the proof of Main Theorem 3 (1). We can also see that the bundle $M_{{\rm B},0}$ is trivial for the structure group. This completes the proof of Main Theorem 3 (2).
@@
By our arguments, we can also check Main Theorem 3 (3) easily. A result close to this is also presented in Theorem 1 (3) of [17]. ”$\overline{D^{\prime}}$” is seen as the ($m+1$)-dimensional bounded and compact manifold $W_{0}$ bounded by $M_{0}$ there. The topology of the ($m+1$)-dimensional manifold is not so precisely discussed there. Only exposition on ”homotopy” is presented there.
@@
Example 1.
Note that $F_{S^{m-n}}(y,t)=t-||y||^{2}$ gives one of simplest cases where $y\in{\mathbb{R}}^{m-n+1}$. For this we can choose $a$ as an arbitrary positive real number. The bundles in (1) and (2) are regarded as ”linear bundles defined before”. Furthermore, in this case, we can consider a suitable situation and regard that ${\Phi}_{F_{S^{m-n}},a}$ is replaced by a diffeomorphism, that the homeomorphism mapping each sphere is regarded as a diffeomorphism, and that (3) can be discussed in the smooth category, for example. In short, we can argue all in the smooth situation in Main Theorem 3. This is also for Main Theorem 2.
This is generalized for more general cases with $m-n\neq 4$ for example. This is thanks to fundamental arguments on singularity theory and differential topology. This is also thanks to well-known classical and sophisticated theory of diffeomorphisms on spheres. This is also related to studies on diffeomorphism groups of unit spheres, studied in [11], and so-called ”pseudo-isotopy”, pioneered in [1], for example.
See also Remark 4, in the next section, for example.
Example 2.
Consider the case of the form $F_{S^{m-n}}(y,t)=t-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ with a Morse function $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$. We have a special generic map $f_{0}$ as in the case of Main Theorem 2.
We also note that Main Theorem 3 respects the case $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y):={\Sigma}_{j=1}^{m-n+1}k_{j,1}{y_{j}}^{2k_{j,2}}$ with $k_{j,1}$ being positive numbers and $k_{j,2}$ being positive integers. See also Remark 3 of [14].
@@
[17] has no essential errors and we have also found several more rigorous arguments. Some of these important arguments are discussed in the present paper. We will revise the version and submit it to the refereed journal, based on comments and our arguments and idea here. Note also that the content of [17] is independent of the content of the present paper essentially.
@@
This new study is based on studies on special generic maps and explicit construction of real algebraic functions and maps of the author. One of related events motivating the author to present this study is the conference ”Extension of the singularity theory” (https://www2.akita-nct.ac.jp/kasedou/workshop/rims2023/index.html)” and related discussions on singularity theory, differential geometry, and real algebraic geometry. See also the previous versions [18].